# Diffusion Denoising as a Certified Defense against Clean-label Poisoning

### v1

@@ 4.3 Comparison to Existing Poisoning Defenses
We finally examine how our defense works better/worse than existing poisoning defenses. We compare ours with six defenses: DP (Ma et al., 2019; Hong et al., 2020), DPA (Levine & Feizi, 2020; Wang et al., 2022), BagFlip (Zhang et al., 2022), ROE (Rezaei et al., 2023), k-NN (Peri et al., 2020), AP (Geiping et al., 2021a), and FrieNDs (Liu et al., 2022). The first four are certified defenses, and the last three are non-certified ones. We test them in the attack settings we use in the previous sections ($\ell_{\infty}$-norm of 16, CIFAR10).
We first compare ours with the filtering-based defenses.
Poisoning attacks TP FP Trained Ours ($\sigma\!=\!0.1$) Witches’ Brew 31 464 ${}^{(86.3)}$   3.0 ${}^{(86.5)}$9.0 Bullseye Polytope 431 63 ${}^{(93.7)}$25.0 ${}^{(93.3)}$0.0
Table 4: Comparison of our defense to Deep kNN.
vs. Deep kNN. Peri et al. (2020) remove poisons from the training data by leveraging the k-nearest-neighbor (kNN) algorithm. They run kNN on the latent representations of the training data to identify potentially malicious samples (i.e., samples with labels different from the nearest ones) and remove them. We compare Deep kNN’s effectiveness to our defense. We use the defense configurations that bring the best results in the original study, i.e., setting $k$ to 500. Table 5 shows our results. Deep kNN fails to remove most poisoning samples (but it removes many benign samples!).
Now we compare ours with the training-time defenses.
Heuristic defense Certified defense Poisoning attacks DP-SGD AT FrieNDs ROE Ours Bullseye Polytope ${}^{(93.9)}$7.0 ${}^{(90.3)}$96.0 ${}^{(90.2)}$10.0 $*{}^{*}$/A ${}^{(93.5)}$   0.0 Witches’ Brew ${}^{(76.0)}$4.0 ${}^{(66.5)}$   2.0 ${}^{(87.6)}$   8.0 ${}^{(70.2)}$10.0 ${}^{(86.5)}$   9.0 Sleeper Agent ${}^{(74.8)}$5.0 ${}^{(69.0)}$   6.0 ${}^{(87.4)}$ 14.0 ${}^{(68.6)}$12.0 ${}^{(86.0)}$17.0 $*{}^{*}$ is incompatible with BP as ROE needs to train 250 models from scratch.
@@
vs. DP-SGD. Ma et al. (2019) proposed a certified defense against poisoning attacks that leverage differential privacy (DP). DP makes a model less sensitive to a single-sample modification to the training data. But the defense offers a weak guarantee in practice; for example, the certificate is only valid in CIFAR10, when an adversary tampers one training sample. (Hong et al., 2020) later empirically shows that DP is still somewhat effective against poisoning. We compare our defense to the training with $(1.0,0.05)$-DP, follow the prior work (Lecuyer et al., 2019), on the same tampered training data. We use Opacus [^2] https://opacus.ai to train models with DP and keep all the other configurations the same. In Table 5, both defenses significantly reduce the attack success to 0–16%. Our defense achieves at most 10% higher accuracy than DP. DP reduces the attack success slightly more against clean-label backdooring.
@@
vs. AT. Geiping et al. (2021a) adapts the adversarial training (AT)for defeating clean-label poisoning: in each mini-batch, instead of crafting adversarial examples, they synthesize poisoning samples and have a model to make correct classifications on them. While effective, this approach could overfit a specific poisoning attack used in the adapted AT, leaving the model vulnerable to unseen poisoning attacks. We thus compare our defense with the original AT with the PGD-7 bounded to $\epsilon$ of 4. Table 5 shows that our defense achieves higher accuracy than the robust models while making the attcaks ineffective. Interestingly, AT cannot defeat the BP attacks. We hypothesize that clean-label poisons in the training data are already adversarial examples; thus, AT may not add more perturbations to them during training.
@@
vs. FrieNDs. Liu et al. (2022) advances the idea of training robust models. Instead of adding random Gaussian noise to the data during training, they use “friendly” noise, pre-computed by a defender, that minimizes the accuracy loss to a model they will train. When their defense is in action, they train a model on the tampered training data with two types of noise: the friendly noise computed before and a weak random noise sampled uniformly within a bound. Ours and FrieNDs greatly reduce the attack success while preserving the model’s utility. The critical difference is that ours is a certified defense, while FrieNDs is not. A promising future work we leave here is developing adaptive clean-label poisoning attacks against FrieNDs.
@@
vs. ROE. Certified defenses (Steinhardt et al., 2017; Ma et al., 2019; Diakonikolas et al., 2019; Levine & Feizi, 2020; Gao et al., 2021; Zhang et al., 2022; Wang et al., 2022; Rezaei et al., 2023) are; however, most of these defenses showed their effectiveness in limited scenarios, such as in MINST-17 (Zhang et al., 2022), a subset of MNIST containing only the samples of the digits 1 and 7. In consequence, they aren’t compatible with our clean-label poisoning settings. A recent defense by Rezaei et al. (2023) demonstrates the certified robustness in practical settings; we thus compare our defense to their run-off-election (ROE). In Table 5, we empirically show that our defense is comparable to ROE while we preserve the model accuracy more.
@@
Note on the computational efficiency. Most training-time defenses require additional mechanisms we need to “add” to the training algorithm. AT needs to craft adversarial examples or clean-label poisons (Geiping et al., 2021a). DP-SGD introduces noise into the gradients during training. FrieNDs pre-computes friendly noise and then optimally applies the noise during the training, and ROE requires training of 250 base classifiers to achieve the improved certificates. In contrast, our defense only requires to run forward once with an off-the-shelf diffusion model.
5 Discussion and Future Work
@@
Our work offers a new perspective on the cat-and-mouse game played in clean-label poisoning.
@@
Our work shows that this cat-and-mouse game need not be as pessimistic as the cat portrays.
@@
Next steps. One of our main conclusions—other than that our defense is effective—is that there is an extremely large gap between the certified accuracy and best practical attack accuracy. When using diffusion models to denoise adversarial examples, a practical adversary can achieve within a few percentage points of the worst-case certified result; for example, if we can certify a lower bound of $50\%$ robust accuracy, a practical adversary could achieve $55\%$ robust accuracy. But in clean-label poisoning, even at perturbation levels where we can’t certify anything, the attacks fail to decrease the model utility at all. This implies either that (1) current attacks are not fully exploiting the space of possible attacks, or (2) that the certificate provided is loose and a more careful analysis could tighten it.
Acknowledgements
This work is supported by Google Faculty Research Award.
@@ Appendix A Experimental Setup in Detail
We adapt the poisoning benchmark from the prior work (Schwarzschild et al., 2021). Most work on clean-label poisoning attacks uses this benchmark for showcasing their attack success. The open-sourced implementation of Sleeper Agent is incompatible with this framework; thus, we run this attack separately. We implement the attacks to use a comparable $\ell_{2}$-norm bound in the poison-crafting process (see §B). We implemented our benchmarking framework in Python v3.7333https://www.python.org/ and PyTorch v1.13444https://pytorch.org. We use the exact attack configurations and training hyper-parameters that the original study employs.
@@ Appendix B Translating $l_{\infty}$-bound into $l_{2}$-bound
Appendix C More Experimental Results
@@
C.1 Using the Evaluation Setup by Schwarzschild et al. (2021)
@@
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Scenario ${}^{\dagger}$0.0 0.5 1.0 1.5 2.0 Poison Frog! (Shafahi et al., 2018) White-box ${}^{(69.8)}$13.0 ${}^{(55.9)}$8.0 ${}^{(43.1)}$4.0 ${}^{(34.7)}$9.0 ${}^{(26.9)}$11.0 Convex Polytope (Zhu et al., 2019) ${}^{(69.8)}$24.0 ${}^{(53.8)}$5.0 ${}^{(42.6)}$9.0 ${}^{(35.0)}$8.0 ${}^{(27.7)}$6.0 Bullseye Polytope (Aghakhani et al., 2021) ${}^{(69.4)}$100.0 ${}^{(55.8)}$10.0 ${}^{(42.8)}$8.0 ${}^{(35.0)}$9.0 ${}^{(27.4)}$9.0 Label-consistent Backdoor (Turner et al., 2019) ${}^{(69.8)}$2.0 ${}^{(55.9)}$3.0 ${}^{(42.7)}$5.0 ${}^{(34.9)}$8.0 ${}^{(27.4)}$12.0 Hidden Trigger Backdoor (Saha et al., 2020) ${}^{(69.8)}$5.0 ${}^{(55.9)}$3.0 ${}^{(42.7)}$10.0 ${}^{(35.2)}$9.0 ${}^{(27.3)}$9.0 Poison Frog! (Shafahi et al., 2018) Black-box ${}^{(67.9)}$7.0 ${}^{(53.8)}$6.0 ${}^{(43.3)}$2.5 ${}^{(35.3)}$8.0 ${}^{(28.8)}$6.5 Convex Polytope (Zhu et al., 2019) ${}^{(67.9)}$4.0 ${}^{(53.7)}$3.0 ${}^{(43.3)}$3.5 ${}^{(35.2)}$3.0 ${}^{(28.8)}$8.0 Bullseye Polytope (Aghakhani et al., 2021) ${}^{(67.7)}$8.0 ${}^{(53.8)}$17.5 ${}^{(43.2)}$16.5 ${}^{(35.3)}$7.5 ${}^{(28.6)}$8.5 Label-consistent Backdoor (Turner et al., 2019) ${}^{(67.9)}$3.5 ${}^{(53.8)}$2.0 ${}^{(43.5)}$4.5 ${}^{(35.2)}$2.5 ${}^{(29.0)}$8.0 Hidden Trigger Backdoor (Saha et al., 2020) ${}^{(67.9)}$7.5 ${}^{(53.7)}$2.0 ${}^{(43.3)}$7.0 ${}^{(35.2)}$10.5 ${}^{(28.7)}$8.5
Table 6: Diffusion denoising against clean-label poisoning (CIFAR10). We denoise the $l_{\infty}$-norm of 8 perturbations added by five poisoning attacks by running a single-step stable diffusion on the entire training set. In each cell, we show the average attack success over 100 runs and the average accuracy of models trained on the denoised data in the parenthesis. Note that ${}^{\dagger}$ indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising.
@@
Results. Diffusion denoising significantly reduces the poisoning attack success. The most successful attack, Bullseye Polytope in the white-box setting, achieves the attack success of 100% in $l_{\infty}$-norm of 8 pixels, but denoising with $\sigma$ of 0.5 can reduce their success to 10%. Our defense reduces the attack success of Poison Frog! and Convex Polytope from 13-34% to 2-8% at $\sigma=0.5$. The two backdoor attacks (label-consistent and hidden trigger) exploiting clean-label poisoning are not successful in the benchmark setup (their success rate ranges from 2-7.5%). We thus could not quantify our defense’s effectiveness against these backdoor attacks. Note that Schwarzschild et al. (2021) also showed these backdoor attacks ineffective, and our finding is consistent with their results. We exclude them from our Tiny ImageNet results in §C.2.
@@
C.2 Tiny ImageNet Results
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Scenario ${}^{\dagger}$0.0 0.1 0.25 0.5 1.0 Poison Frog! (Shafahi et al., 2018) WB ${}^{(58.9)}$79.0 ${}^{(56.3)}$3.0 ${}^{(53.3)}$2.0 ${}^{(46.5)}$2.0 ${}^{(32.0)}$1.0 Convex Polytope (Zhu et al., 2019) ${}^{(58.9)}$95.0 ${}^{(56.3)}$14.6 - - - Bullseye Polytope (Aghakhani et al., 2021) ${}^{(58.8)}$100.0 ${}^{(57.4)}$100.0 - ${}^{(48.7)}$0.0 - Poison Frog! (Shafahi et al., 2018) BB ${}^{(58.0)}$5.0 ${}^{(51.9)}$0.0 ${}^{(46.5)}$0.0 ${}^{(36.0)}$1.0 ${}^{(22.6)}$1.0 Convex Polytope (Zhu et al., 2019) ${}^{(58.1)}$0.0 ${}^{(51.9)}$0.0 - - - Bullseye Polytope (Aghakhani et al., 2021) ${}^{(57.9)}$9.0 ${}^{(59.1)}$33.3 - ${}^{(49.8)}$16.7 -
Table 7: Diffusion denoising against clean-label poisoning (Tiny ImageNet). We consider four attacks with the $\ell_{\infty}$-norm of 16 perturbation bound. In each cell, the attack success and accuracy of models trained on the denoised data in the parenthesis on average over 100 runs. Note that ${}^{\dagger}$ indicates the runs with $\sigma$=0.0, representing no-defense scenario.

### v2

@@ 4.3 Comparison to Existing Poisoning Defenses
We finally examine how our defense works better/worse than five existing defenses: k-NN (Peri et al., 2020), DP-SGD (Ma et al., 2019; Hong et al., 2020), AT (Geiping et al., 2021a), FrieNDs (Liu et al., 2022), and ROE (Rezaei et al., 2023). The ROE is a certified defense, and the other four are non-certified ones. We use $\ell_{\infty}$-norm of 16 in CIFAR10.
Table 4: Comparison of ours to Deep kNN.
Poisoning attacks TP FP Trained Ours ($\sigma\!=\!0.1$) Witches’ Brew 31 464 (86.3)   3.0 (86.5)9.0 Bullseye Polytope 431 63 (93.7)25.0 (93.3)0.0
Deep kNN removes poisons from the training data by leveraging the k-nearest-neighbor (kNN) algorithm. They run kNN on the latent representations of the training data to identify potentially malicious samples (i.e., samples with labels different from the nearest ones) and remove them. We compare Deep kNN’s effectiveness to our defense. We use the defense configurations that bring the best results in the original study, i.e., setting $k$ to 500. Table 5 shows our results. Deep kNN fails to remove most poisoning samples (but it removes many benign samples!).
@@
Heuristic defense Certified defense Poisoning attacks DP-SGD AT FrieNDs ROE Ours Bullseye Polytope (93.9)7.0 (90.3)96.0 (90.2)10.0 N/A (93.5)   0.0 Witches’ Brew (76.0)4.0 (66.5)   2.0 (87.6)   8.0 (70.2)10.0 (86.5)   9.0 Sleeper Agent (74.8)5.0 (69.0)   6.0 (87.4) 14.0 (68.6)12.0 (86.0)17.0 ROE is incompatible with BP as ROE needs to train 250 models from scratch.
@@
DP-SGD. Ma et al. (2019) proposed a certified defense against poisoning attacks that leverage differential privacy (DP). DP makes a model less sensitive to a single-sample modification to the training data. But the guarantee is weak in practice; for example, the certificate is only valid in CIFAR10, when an adversary tampers one training sample. Hong et al. (2020) later empirically shows that DP is still somewhat effective against poisoning. We compare our defense to the training with $(1.0,0.05)$-DP, follow the prior work (Lecuyer et al., 2019). In Table 5, DP significantly reduces the attack success to 0–16%. Our defense achieves at most 10% higher accuracy than DP. DP reduces the attack success slightly more against clean-label backdooring.
@@
AT. Geiping et al. (2021a) adapts the adversarial training (AT) for defeating clean-label poisoning: in each mini-batch, instead of crafting adversarial examples, they synthesize poisoning samples and have a model to make correct classifications on them. While effective, the model could overfit a specific poisoning attack used in the adapted AT, leaving the model vulnerable to unseen poisoning attacks. We thus compare our defense with the original AT with the PGD-7 and $\varepsilon$ of 4, assuming that clean-label poisons are already adversarial examples. AT may not add more perturbations to them during training. In Table 5, our defense achieves higher accuracy than the robust models while making the attacks ineffective. Interestingly, AT cannot defeat the BP.
@@
FrieNDs. Liu et al. (2022) advances the idea of training robust models. Instead of adding random Gaussian noise to the data during training, they use “friendly” noise, pre-computed by a defender, that minimizes the accuracy loss to a model they will train. When their defense is in action, they train a model on the tampered training data with two types of noise: the friendly noise computed before and a weak random noise sampled uniformly within a bound. Ours and FrieNDs greatly reduce the attack success while preserving the model’s utility. The critical difference is that ours is a certified defense, while FrieNDs is not. A future work we leave is developing adaptive attacks against FrieNDs.
@@
ROE. Certified defenses (Steinhardt et al., 2017; Ma et al., 2019; Diakonikolas et al., 2019; Levine & Feizi, 2020; Gao et al., 2021; Zhang et al., 2022; Wang et al., 2022; Rezaei et al., 2023) are; however, most of these defenses showed their effectiveness in limited scenarios, such as in MINST-17 (Zhang et al., 2022), a subset of MNIST containing only the samples of the digits 1 and 7. In consequence, they aren’t compatible with our clean-label poisoning settings. A recent defense by Rezaei et al. (2023) demonstrates the certified robustness in practical settings; we thus compare our defense to their run-off-election (ROE). In Table 5, we empirically show that our defense is comparable to ROE while we preserve the model accuracy more.
@@
Note on the computational efficiency. To be practically effective, most training-time defenses require additional mechanisms we need to “add” to the training algorithm. AT craft adversarial examples or clean-label poisons (Geiping et al., 2021a). DP-SGD introduces noise into the gradients during training. FrieNDs pre-computes friendly noise and then optimally applies the noise during the training, and ROE requires training of 250 base classifiers to achieve the improved accuracy. In contrast, we only require to run forwards once with an off-the-shelf diffusion model.
5 Discussion
@@
Our work offers a new perspective on the cat-and-mouse game played in clean-label poisoning research.
@@
This game need not be as pessimistic as the cat portrays.
@@ Appendix A Experimental Setup in Detail
We adapt the poisoning benchmark from the prior work (Schwarzschild et al., 2021). Most work on clean-label poisoning attacks uses this benchmark for showcasing their attack success. The open-sourced implementation of Sleeper Agent is incompatible with this framework; thus, we run this attack separately. We implement the attacks to use a comparable $\ell_{2}$-norm bound in the poison-crafting process (see §B). We implemented our benchmarking framework in Python v3.7111https://www.python.org/ and PyTorch v1.13222https://pytorch.org. We use the exact attack configurations and training hyper-parameters that the original study employs.
@@ Appendix B Translating $l_{\infty}$-bound into $l_{2}$-bound
Appendix C Proof for Our Provable Guarantee
This section provides the proof of our provable robustness against clean-label poisoning attacks within the $\ell_{2}$-norm.
Studies on certified robustness. We first note that certified robustness has been explored for other types of attacks, such as adversarial examples (Cohen et al., 2019; Salman et al., 2020; Carlini et al., 2022) and backdoor attacks (Weber et al., 2023), which assumes different threat models compared to clean-label poisoning. Cohen et al. (2019) (certified defenses against adversarial examples) operate under the assumption that the target classifier remains stationary, whereas clean-label poisoning involves the victim “trains” the classifier on the compromised data. Weber et al. (2023) presents the provable guarantee against backdooring, where classifiers are trained on datasets containing backdoor poisons. But, backdoor attacks differ substantially as they assume that the adversary can manipulate test-time inputs with specific trigger patterns.
Comparison to the prior studies. Our provable guarantee can be viewed as a special case of the backdoor robustness framework proposed by Weber et al. (2023), where the backdoor adversary applies zero perturbations ($\varepsilon\!=\!0$) to test-time inputs. We also highlight that our use of diffusion denoising as a randomized smoothing mechanism enhances classification accuracy by 1.8$\times$ compared to the results reported by (Weber et al., 2023). Moreover, to achieve the same level of certified accuracy, we need 20$\times$ less computational cost compared to prior studies (Cohen et al., 2019; Weber et al., 2023).
Proof. On a high-level, we first review the robustness guarantee provided by (Cohen et al., 2019) to adversarial examples, specifically for $\ell_{2}$-norm perturbations in the input space. Next, we elaborate on how Weber et al. (2023) extend the previous robustness—achieved through training with randomized smoothing. By smoothing the training data first and then training a classifier on it, Weber et al. (2023) establish a provable guarantee against backdooring—an attack exploits data poisoning. We finally reduce the robustness guarantee against backdooring to a special case where no input perturbations occur at test-time, thereby providing a provable guarantee against clean-label poisoning.
Theorem C.1 (Robustness guarantee by Cohen et al. (2019)).
Let $f:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ be any deterministic (or random) function, $\varepsilon\sim\mathcal{N}(0,\sigma^{2},I)$, and $g$ to be the smoothed classifier robust under $\varepsilon$. Suppose $c_{A}\in\mathcal{Y}$ and $p_{A},p_{B}\in[0,1]$ satisfy:
$\mathbb{P}(f(x+\varepsilon)=c_{A})\geq p_{A}>p_{B}\geq\underset{c\neq c_{A}}{\max}\mathbb{P}(f(x+\varepsilon)=c)$
Then $g(x+\delta)=c_{A}$ for all $||\delta||_{2}<R$, where
$R=\frac{\sigma}{2}\big{(}\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)}$
Theorem C.1 establishes that the smoothed classifier $g$ is robust around a test-time input $x$ within the $\ell_{2}$-norm radius of $R$ as specified above. Here $\Phi$ denotes the inverse of the standard Gaussian CDF. For the proof of this theorem, we refer readers to Appendix A of the original study by Cohen et al. (2019).
The previous approach by Cohen et al. (2019) does not impose specific restrictions on how the smoothed classifier $g$ is constructed, A common method they employ is to apply additive Gaussian noise to the training data at each mini-batch during training. However, this method is not compatible with our setting where we train a classifier on the smoothed training data. To address this challenge, we adopt the approach proposed by Weber et al. (2023), which relies on a smoothed classifier $g$ obtained through a method that aligns with our framework, defined as follows:
Definition C.2 (Smoothed classifier considered by Weber et al. (2023)).
Suppose a base classifier $h$, defined as:
$h(x,\mathcal{D})=\operatorname*{arg\,max}_{y}p(y|x,\mathcal{D}),$
where $p$ is learned from the training data $\mathcal{D}$ and to maximize the conditional probability distribution over labels $y$. Then the smoothed classifier $q$ is defined by:
$q(y|x,\mathcal{D})=\mathbb{P}_{X,D}(h(x+X,\mathcal{D}+D)=y),$
where $X\sim\mathbb{P}_{X}$ represents random variables, and $D\sim\mathbb{P}_{D}$ denotes smoothing distributions. $D$ is a collection of $n$ i.i.d random variables, composed of $D^{i}$s, where each $D^{i}$ will be added to a training samples in $\mathcal{D}$. $\mathbb{P}_{X}$ is the noise distribution for test-time inputs, and $\mathbb{P}_{D}$ is the noise distribution for the training data. The final, smoothed classifier then assigns the most likely class to an instance $x$ under $q$, so that:
$g(x,\mathcal{D})=\operatorname*{arg\,max}_{y}q(y|x,\mathcal{D})$
Note that under this formulation, the smoothed classifier can defend against adversarial examples by setting the training set noise to zero, $D\equiv 0$. Similarly, setting the test-time noise to zero, $X\equiv 0$, enables the classifier to defend against clean-label poisoning. We will first follow the robustness framework by Weber et al. (2023) and then show the special case, $X\equiv 0$, establishing our guarantee against clean-label poisoning.
Theorem C.3 (Robustness guarantee by Weber et al. (2023)).
Suppose $q$ is the smoothed classifier defined in Def. C.2 with smoothing distribution $Z:=(X,D)$ with $X$ taking values in $\mathbb{R}^{d}$ and $D$ being a collection of $n$ independent $\mathbb{R}^{d}$-valued random variables, $D=(D^{1},D^{2},...,D^{n})$. Let $\Omega_{x}\in\mathbb{R}^{d}$ and $\Delta:=(\delta^{1},\delta^{2},...,\delta^{n})$ for backdoor patterns $\delta^{i}\in\mathbb{R}^{d}$. Let $y_{A}\in C$ and $p_{A},p_{B}\in[0,1]$ such that $y_{A}=g(x,D)$ and
$q(y_{A}|x,\mathcal{D})\geq p_{A}>p_{B}\geq\underset{y\neq y_{A}}{\max}q(y|x,\mathcal{D}).$
If the optimal type II errors, for testing the null $Z\sim\mathbb{P}_{0}$ against the alternative $Z+(\Omega_{x},\Delta)\sim\mathbb{P}_{1}$, satisfy
$\beta^{*}(1-p_{Z};\mathbb{P}_{0},\mathbb{P}_{1})+\beta^{*}(p_{B};\mathbb{P}_{0},\mathbb{P}_{1})>1,$
then it is guaranteed that $y_{A}=\operatorname*{arg\,max}_{y}q(y|x+\Omega_{x},\mathcal{D}+\Delta)$.
Theorem C.3 by Weber et al. (2023) establishes that the smoothed classifier $q$, as defined in Definition C.2, trained on data containing backdoor patterns $\Delta$, is robust around a test-time input $x$ with the perturbations in $\Omega_{x}$. For the proof of this theorem, we refer readers to Appendix A.1 of the original study by Weber et al. (2023).
Corollary C.4 ($\ell_{2}$-robustness achieved by Gaussian smoothing in Weber et al. (2023)).
Let $\Delta=(\delta_{1},\delta_{2},...,\delta_{n}$), $\Omega_{x}$ be $\mathbb{R}^{d}$-valued backdoor patterns, and $\mathcal{D}$ be a training dataset. Suppose that for each $i$, the smoothing noise on the training features is $D^{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\sigma^{2}\mathbbm{1}_{d})$. Let $y_{A}\in C$ such that $y_{A}=q(x+\Omega_{x},\mathcal{D}+\Delta)$ with class probabilities satisfying:
$q(y_{A}|x+\Omega_{x},\mathcal{D}+\Delta)\geq p_{A}>p_{B}\geq\underset{y\neq y_{A}}{\max}q(y|x+\Omega_{x},\mathcal{D}+\Delta)$
Then if the backdoor patterns on the training samples are bounded by
$\sqrt{\sum_{i=1}^{n}||\delta_{i}||_{2}^{2}}<\frac{\sigma}{2}\big{(}\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)},$
it is guaranteed $y_{A}=q(x+\Omega_{x},\mathcal{D})=q(x+\Omega_{x},\mathcal{D}+\Delta)$.
Now we can simplify Corollary C.4 by assuming an adversary compromises $r\leq n$ training samples with the largest noise pattern’s $\ell_{2}$-norm bound, $||\delta||_{2}$:
$||\delta||_{2}<\frac{\sigma}{2\sqrt{r}}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)},$
A typical choice of $r$ in clean-label poisoning is $\sim$10of the training data. But when computing our certificate bound in §3, we consider the worst-case adversary capable of compromising the entire training set. In this case, $r\!=\!n$, which becomes:
$||\delta||_{2}<\frac{\sigma}{2\sqrt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n}}}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)}\leq\frac{\sigma}{2\sqrt{r}}(\Phi^{-1}(p_{A})-\Phi^{-1}(p_{B})\big{)},$
Under this formulation, we observe that against the strongest adversary who can compromise the entire training data $n$, the robustness bound is much smaller compared to that against the realistic attacker capable of compromising only 10% of the training data.
Certifying the robustness with multiple smoothed classifiers. In our certification process, a defender trains $n$ models [^3] Note that this $n$ is different from $n$ we use for denoting the cardinality of the training set. that return the prediction of $x$, the target test-time sample, as $y_{A}$. Then the lower bound of $p_{A}$ becomes $\alpha^{1/n}$, with the probability of at least $1-\alpha$. Plugging this into the above equation yields:
$||\delta||_{2}<\frac{\sigma}{\sqrt{n}}(\Phi^{-1}(p_{A})\big{)}$
Corollary C.5 (Our robustness guarantee against clean-label poisoning.).
Consider the smoothed classifier $q$ from Theorem C.3, with no test-time input perturbation $X$ and the smoothing mechanism being diffusion denoising, which applies a Gaussian process to add and remove noise from the input. Under this setting, Theorem C.3 provides the same robustness guarantee for clean-label poisoning attacks.
Proof.
By setting $X\equiv 0$ in Theorem C.3 (and therefore $\Omega_{x}$), the smoothed classifier will output the same label prediction $y_{A}$ within the same $\ell_{2}$-bound for the perturbations applied to the training data of size $n$. ∎
We also note that Weber et al. (2023) formally demonstrate how to derive a robustness guarantee for $\ell_{\infty}$-norm perturbations using uniform smoothing. Because our work employs off-the-shelf diffusion denoising models that rely on Gaussian random noise, we do not provide a provable guarantee for $\ell_{\infty}$-based clean-label poisoning. (but we do empirically show substantial effectiveness against $\ell_{\infty}$-based attacks!) We believe that extending our work to provide formal guarantees for $\ell_{\infty}$-based attacks would be straightforward and thus leave this for future work.
Appendix D More Experimental Results
@@
D.1 Using the Evaluation Setup by Schwarzschild et al. (2021)
@@
Table 6: Diffusion denoising against clean-label poisoning (CIFAR10). We denoise the $l_{\infty}$-norm of 8 perturbations added by five poisoning attacks by running a single-step stable diffusion on the entire training set. In each cell, we show the average attack success over 100 runs and the average accuracy of models trained on the denoised data in the parenthesis. Note that indicates the runs with $\sigma$=0.0, the same as our baseline that trains models without any denoising.
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Scenario 0.0 0.5 1.0 1.5 2.0 Poison Frog! (Shafahi et al., 2018) White-box (69.8)13.0 (55.9)8.0 (43.1)4.0 (34.7)9.0 (26.9)11.0 Convex Polytope (Zhu et al., 2019) (69.8)24.0 (53.8)5.0 (42.6)9.0 (35.0)8.0 (27.7)6.0 Bullseye Polytope (Aghakhani et al., 2021) (69.4)100.0 (55.8)10.0 (42.8)8.0 (35.0)9.0 (27.4)9.0 Label-consistent Backdoor (Turner et al., 2019) (69.8)2.0 (55.9)3.0 (42.7)5.0 (34.9)8.0 (27.4)12.0 Hidden Trigger Backdoor (Saha et al., 2020) (69.8)5.0 (55.9)3.0 (42.7)10.0 (35.2)9.0 (27.3)9.0 Poison Frog! (Shafahi et al., 2018) Black-box (67.9)7.0 (53.8)6.0 (43.3)2.5 (35.3)8.0 (28.8)6.5 Convex Polytope (Zhu et al., 2019) (67.9)4.0 (53.7)3.0 (43.3)3.5 (35.2)3.0 (28.8)8.0 Bullseye Polytope (Aghakhani et al., 2021) (67.7)8.0 (53.8)17.5 (43.2)16.5 (35.3)7.5 (28.6)8.5 Label-consistent Backdoor (Turner et al., 2019) (67.9)3.5 (53.8)2.0 (43.5)4.5 (35.2)2.5 (29.0)8.0 Hidden Trigger Backdoor (Saha et al., 2020) (67.9)7.5 (53.7)2.0 (43.3)7.0 (35.2)10.5 (28.7)8.5
@@
Results. Diffusion denoising significantly reduces the poisoning attack success. The most successful attack, Bullseye Polytope in the white-box setting, achieves the attack success of 100% in $l_{\infty}$-norm of 8 pixels, but denoising with $\sigma$ of 0.5 can reduce their success to 10%. Our defense reduces the attack success of Poison Frog! and Convex Polytope from 13-34% to 2-8% at $\sigma=0.5$. The two backdoor attacks (label-consistent and hidden trigger) exploiting clean-label poisoning are not successful in the benchmark setup (their success rate ranges from 2-7.5%). We thus could not quantify our defense’s effectiveness against these backdoor attacks. Note that Schwarzschild et al. (2021) also showed these backdoor attacks ineffective, and our finding is consistent with their results. We exclude them from our Tiny ImageNet results in §D.2.
@@
D.2 Tiny ImageNet Results
Table 7: Diffusion denoising against clean-label poisoning (Tiny ImageNet). We consider four attacks with the $\ell_{\infty}$-norm of 16 perturbation bound. In each cell, the attack success and accuracy of models trained on the denoised data in the parenthesis on average over 100 runs. Note that indicates the runs with $\sigma$=0.0, representing no-defense scenario.
Our defense against $\ell_{\infty}$ attacks at $\sigma$ (%) Poisoning attacks Scenario 0.0 0.1 0.25 0.5 1.0 Poison Frog! (Shafahi et al., 2018) WB (58.9)79.0 (56.3)3.0 (53.3)2.0 (46.5)2.0 (32.0)1.0 Convex Polytope (Zhu et al., 2019) (58.9)95.0 (56.3)14.6 - - - Bullseye Polytope (Aghakhani et al., 2021) (58.8)100.0 (57.4)100.0 - (48.7)0.0 - Poison Frog! (Shafahi et al., 2018) BB (58.0)5.0 (51.9)0.0 (46.5)0.0 (36.0)1.0 (22.6)1.0 Convex Polytope (Zhu et al., 2019) (58.1)0.0 (51.9)0.0 - - - Bullseye Polytope (Aghakhani et al., 2021) (57.9)9.0 (59.1)33.3 - (49.8)16.7 -